OpenStudy (anonymous):

@zepdrix Please help me! How long does it take $1075 to double if it is invested at 4% interest compounded quarterly? a. 19.7 years b.19.7 months c. 17.4 years d. none of these

zepdrix (zepdrix):

\[\Large\bf\sf A\quad=\quad P\left(1+\frac{r}{n}\right)^{nt}\]

zepdrix (zepdrix):

\[\Large\bf\sf r=0.04\]\[\Large\bf\sf n=4\]Quarterly means 4 times per year. \[\Large\bf\sf P=1075\]

zepdrix (zepdrix):

How long does it take to double. So in other words, when does A = twice P.\[\Large\bf\sf 2150\quad=\quad 1075\left(1+\frac{0.04}{4}\right)^{4t}\]

zepdrix (zepdrix):

Do you understand the setup?

zepdrix (zepdrix):

Btw, cats rule -_- dogs drool. end of story.

OpenStudy (anonymous):

Totaaly get it! =) Thanks!

zepdrix (zepdrix):

Understand how to solve for t? Gotta use logs and stuff :o

zepdrix (zepdrix):

Or you could plug in your multiple choice options I suppose :3

OpenStudy (anonymous):

I got it! Thank you so much! I think the answer is 17.3!

OpenStudy (anonymous):

I mean 17.4 thanks for everything!